OpenStudy (anonymous):

Solve for the indicated variable 3x-4y=12; Y

hero (hero):

y = (12 - 3x)/-4 = -3 + 3x/4 = 3x/4 - 3

OpenStudy (anonymous):

I don't understand this

OpenStudy (campbell_st):

the question is asking you to make y the subject. 3x - 4y = 12 or 3x - 12 = 4y divide by 4 (3x-12)/4 = y

hero (hero):

You're solving for the variable y. In other words, you're manipulating the equation in a way such that you isolate the variable y on one side of the equation, and write the rest of the terms on the other side of the equation in a simple form.

OpenStudy (anonymous):

okay i think i understand now..so all you haev to do is just find the answer for y?